# Assortment Optimization Under History-Dependent Effects

### v2

@@ 5.2 Assortment Planning
In this section, we use the formulation proposed in Section 3 to optimize weekly menus using the above estimators. We design weekly workday menus because the cafeteria does not frequently operate on weekends, and in practice, the cafeteria indeed plans for a week. We randomly select a month [^8] Our estimation includes monthly fixed effects. Hence, the product’s utility will vary month by month. and use the estimated parameters in Table 3 as the input for our conic models. To keep the menu sizes comparable, we impose a daily cardinality constraint of $6$, which is the median assortment size in the dataset.
Figure 2: Offering times, price, and base attraction value of offered dishes
@@ 5.2.1 Assortment Patterns
Figure 2 shows the total number of products offered, the prices, and the base attraction values obtained for the assortments under the MNL model, under history-dependent MNL models with different memory lengths, and under an assortment planning used in current practice. The current practice assortment planning was randomly chosen from data on the same month which was used for computing the base attraction value when optimization (the base utility contains month fixed effects).
As shown in Figure 2 (a) and (b), the MNL-based model tends to offer fewer products with higher prices. Moreover, the number of products that are offered increases with the memory length $M$, because it is beneficial to offer more products and avoid repeating the same product within the memory length when satiation effects last for a long period. When the memory length increases, the assortments include more products with lower base attraction values, as shown in Figure 2 (c).
Figure 3: Assortment details
Figure 3 presents assortment details for examining how the memory length affects assortment patterns. Figure 3 (a)-(d) show the assortments obtained from MNL-based models with $M=0,1,2,3$. The horizontal axes represent products indexed from the highest to the lowest price. The vertical axes represent the days of a week. The products offered are colored blue. The subfigures show that the interval between two adjacent offerings of the same dish grows when $M$ increases, and all optimal assortments represent a cyclic-style structure. The interval for two adjacent offerings of most products is just the memory length $M$, which suggests that the cycle length is $M+1$, aligning with the $(M+1)$-cyclic policy in Section 4. Moreover, it can be observed that products with higher prices or lower base attraction values are more sensitive to the satiation effect, and thus will be offered at most once within the span of the memory length. For instance, the high-price products 1 to 4 strictly repeat in intervals of $M$ days to avoid the satiation effect, possibly because the revenue loss caused by the utility deterioration will be scaled up by the high price. On the other hand, products 6 and 12 are provided on two consecutive days when $M=2$ due to their relatively high base attraction value. The price and base attraction value for each product are shown in Figure 3 (e) and (f) for reference.
@@ 5.2.2 Performance Comparison
In this section, we focus on two metrics to evaluate the assortment performance: variety and revenue, as these are concerns for the corporation cafeteria. We adopt the Herfindahl-Hirschman Index (HHI), a measure of market concentration (Rhoades 1993), to quantify the variety of an assortment planning. The HHI value is $\sum_{i\in[n]}(\frac{h_{i}}{H})^{2}$, where $n$ is the total number of products, $h_{i}$ is the number of times product $i$ is offered, and $H=\sum_{i\in[n]}h_{i}$. A low HHI value reflects a large variety. To compute the out-of-sample revenue, we set the model with the best prediction performance for the second data set, i.e., the data set in the fourth quarter, as the “ground truth” model. Besides the assortments discussed in Section 5.2.1, we additionally compute assortments obtained by the $(M+1)$-cyclic policy with non-overlapping constraints that is presented in Section 4.
Figure 4: Assortment variety and revenue
Figure 4 illustrates the relationship between the variety and revenue of different assortments. The horizontal axis represents the HHI value, and the vertical axis represents the normalized out-of-sample average daily revenue. We normalize the out-of-sample revenue of different assortment plannings with the revenue of current practice as a benchmark. Generally, assortments with a low HHI value can help avoid the negative impacts of satiation effects and maintain high product utility. Figure 4 shows that the assortments generated by our model ($M>0$) have a lower HHI and higher revenue than the assortments obtained with the MNL model. For instance, the average revenue obtained using our model with $M=2$ is approximately $19\%$ greater than the revenue under the MNL model. The $(M+1)$-cyclic policy also performs well. With the same memory length $M$, the $(M+1)$-cyclic policy increases the assortment variety via the non-overlapping constraint and even generates more revenue than the assortments without non-overlapping constraints. We also note that, in practice, assortments that are determined manually have low HHI values but carry substantial revenue losses.
In conclusion, our case study first calibrates the history-dependent effect using transaction data from a corporation cafeteria. Then we demonstrate that offering a wide variety of products with a low frequency can efficiently compensate for the satiation effect. Finally, our findings highlight the advantages of our model, which captures the complex trade-off between products’ utility, their revenue, and history-dependent effects.
@@ 6 Numerical Study
This section presents the results of our numerical studies on synthetic data. In Section 6.1, we illustrate the efficiency of formulation (Conic) in terms of computation time and solution quality. In Section 6.2, we demonstrate the performance of cyclic policies. All our numerical studies were implemented in the Julia, version 1.8.2 (Bezanson et al. 2017), using modeling language JuMP (Lubin et al. 2023), on a Virtual Machine with 64GB RAM and a 32-core Intel Xeon (Skylake, IBRS) @2.39 GHz processor. All mixed-integer linear programs were solved using Gurobi 11.0.0 (Gurobi Optimization, LLC 2023) and mixed-integer exponentical cone programs were solved using Mosek 10.2 (MOSEK 2022).
@@ 6.1 Formulation Efficiency
We consider four alternate formulations for solving (HAP):
@@
  • $\mathsf{Env}$: This is the mixed-integer linear formulation obtained by replacing the exponential cone constraint in (Conic) with the convex envelope based constraints (8), as detailed in Remark 3.4.
  • $\mathsf{ML}$: This is the mixed-integer linear formulation based on multilinear extension of attraction value functions, and its deviation is detailed in Appendix 11.1.
  • $\mathsf{SCIP}$: This refers to solving the binary nonlinear formulation (HAP) using the global optimization solver SCIP.
We compare the performance of the four formulations on synthetically generated problem instances. We generate these instances as follows. For a fixed number of products $N$ and memory length $M$, the base utility of each product $\beta_{i}^{0}$ is drawn from a uniform $[-1,1]$ distribution and its revenue $r_{i}$ from a uniform $[1,10]$ distribution. The no-purchasee utility is 00. For each $i\in[N]$ and $m\in[M]$, we set the history-dependent effect $\beta_{i}^{m}$ from two uniform distributions $U[-1,0]$ and $U[-2,-1]$ to represent a weak and strong satiation effect, denoted as $W$ and $S$, respectively. We vary $N\in\{30,50,70\}$, $M\in\{1,2\}$, and $T\in\{5,10\}$. For each combination of $N$, $M$, and a satiation level (W or S), we generate five instances of $\{r_{i},\beta_{i}^{0},\beta_{i}^{m}\}_{i\in[N],m\in[M]}$.
We solve each instance of (HAP) using one of the formulations $\mathsf{Conic}$, $\mathsf{Env}$, $\mathsf{ML}$, and $\mathsf{SCIP}$. We report the computation time and the end gap, which is defined as
$G_{\textsf{end}}=100\%\times\frac{R_{\mathsf{U}}-R_{\mathsf{IP}}}{R_{\mathsf{IP}}},$
where $R_{\mathsf{U}}$ and $R_{\mathsf{IP}}$ is the best upper bound and the best integer solution at termination with a time limit of 3600s and an optimality gap tolerance of $0.5\%$, respectively.
Table 4: Formulation efficiency comparison
@@
Besides the computation time and end gap, it is also helpful to compare the continuous relaxation of different formulations. We compute the root gap as follows
$G_{\textsf{root}}=100\%\times\frac{R_{\textsf{Rlx}}-R_{\textsf{IP}}}{R_{\textsf{IP}}},$
where $R_{\mathsf{Rlx}}$ is the objective value of the continuous relaxation of a given formulation. A small root gap indicates a tight formulation. Table 5 reports the average root gap of Conic, Env, and ML under different parameter configurations. Overall, the root gap of our formulations Conic and Env is significantly smaller than that of ML. It indicates the tightness of our formulations. Moreover, the root gap of our formulations is stable and does not scale with the problem size. For instance, the root gap of Env is no more than $6\%$ when $M=1$ and no more than $16\%$ when $M=2$.
@@
5 1 30 1 3.86 3.22 156.73 - 10 1 30 1 4.38 3.65 179.23 -
5 1 30 2 6.81 5.89 301.08 - 10 1 30 2 7.83 6.56 343.77 -
5 1 50 1 2.89 2.5 295.99 - 10 1 50 1 3.3 2.82 336.43 -
5 1 50 2 5.64 5.04 517.19 - 10 1 50 2 7.11 5.64 588.26 -
5 1 70 1 3.95 3.55 400.53 - 10 1 70 1 4.61 4.77 455.25 -
5 1 70 2 3.84 3.45 713.79 - 10 1 70 2 5.05 4.2 810.67 -
5 2 30 1 4.57 4.58 224.85 - 10 2 30 1 5.64 5.61 275.29 -
5 2 30 2 13.17 13.17 587.57 - 10 2 30 2 15.4 15.32 786.57 -
5 2 50 1 4.77 4.77 430.75 - 10 2 50 1 5.9 5.76 539.85 -
5 2 50 2 10.1 10.06 921.6 - 10 2 50 2 13.34 12.09 1419.53 -
5 2 70 1 5.47 5.3 717.98 - 10 2 70 1 6.94 6.46 940.42 -
5 2 70 2 6.85 6.43 1243.68 - 10 2 70 2 9.02 7.85 1890.09 -
6.2 Computing Two-Cyclic Policies
@@
In this section, we demonstrate the efficiency and scalability of (11) for computing two-cyclic policies. We propose a projected cutting plane algorithm to implement (11). Note that due to the presence of $\boldsymbol{\Gamma}$ variables, the number of variables in (11) is $O(N^{2})$. To address this issue, we propose a projected cutting plane algorithm to generate a compact approximation of (11). The cutting plane algorithm first sets the linearized (2-Cyclic) as a base formulation and obtains a fractional solution using the continuous relaxation of the base formulation. The cutting plane algorithm then uses a separation oracle to generate new cuts, which are the projection of constraints (11c)-(11e) violated by the fractional solution to the space of $(\boldsymbol{x},\boldsymbol{\gamma})$ variables. Adding new cuts into the base model leads to a formulation with a tighter continuous relaxation. Running the separation oracle once is a round of cut generation. We can add new cuts for $K$ rounds and the final formulation as BF-$K$, where BF denotes “bound-free”. Indeed, one can repeat this process (i.e., a large $K$) until the fractional solution of BF-$K$ does not violate (11c)-(11e). We observe that the marginal improvement of formulation tightness is decreasing. Hence, we can set a small $K$ and obtain a much tighter formulation than the base model. More specifically, we set $K=1$ to compute the two-cyclic policy because BF-$1$ has the most marginal improvement of the root gap. Recall that a small root gap means a tight continuous relaxation. The relative improvement in the root gap of BF-$K$ compared to the conic formulation (Cycle-Conic) with non-overlapping constraint is
$G_{\textsf{root}}^{K}=100\%\times\frac{G_{\textsf{root,conic}}-G_{\textsf{root,BF-K}}}{G_{\textsf{root,conic}}},$
where $G_{\textsf{root,BF-K}}$ and $G_{\textsf{root,conic}}$ are the root gap of BF-$K$ and (Cycle-Conic), respectively.
@@
Table 6: The computational performance of (11) with a cutting plane implementation
N $\boldsymbol{\beta}$ $G^{1}_{\mathsf{root}}(\%)$ $G^{2}_{\mathsf{root}}(\%)$ $G^{3}_{\mathsf{root}}(\%)$ $T_{\mathsf{BF}-1}$(s)
500 W 98.36 98.54 98.66 426
500 S 98.53 98.71 98.80 523
750 W 98.67 98.83 98.90 579
750 S 98.57 98.74 98.83 604
1000 W 98.79 98.89 98.96 606
1000 S 98.90 99.01 99.09 611
Table 6 summarizes the performance of different formulations. BF-$1$ has the most marginal improvement of the root gap. After adding one-round cuts, the natural continuous relaxation of BF-$1$ approximately shrinks $98\%$ root gap of the conic formulation (Cycle-Conic). It suggests that the natural continuous relaxation of BF-$1$ is extremely close to the optimal integer solution and much tighter than the continuous relaxation of (Cycle-Conic), which supports our theoretical result in Theorem 4.4. The last column reports the average running time, including the time for cut-generation, of using formulation BF-$1$ to compute an optimal two-cyclic policy. The average computation time for large instances (e.g., $N=1000$) is about $600$ seconds. This shows the computational efficiency and scalability of the bound-free formulation (11).

### v3

@@ 5.2 Assortment Planning
In this section, we use the formulation proposed in Section 3.1 to optimize weekly menus using the above estimators. We design weekly workday menus because the cafeteria does not frequently operate on weekends, and in practice, the cafeteria indeed plans for a week. We use the estimated parameters in Table 5.1 as the input for our conic models. To keep the menu sizes comparable, we impose a daily cardinality constraint of $6$, which is the median assortment size in the dataset.
@@ 5.2.1 Assortment Patterns
Figure 5.2.1 shows the accumulated number of products for different offering times, the prices, and the base attraction values obtained from the assortments under the MNL model, under history-dependent MNL models with different memory lengths, and under an assortment planning used in practice. The current practice assortment was randomly chosen from the second test data set.
[…]
(a) Offering times
(b) Price
(c) Base attraction value
Information of products offered in assortments obtained by different models
As shown in Figure 2(a) and 2(b), the MNL-based model tends to offer fewer products with higher prices. Moreover, the number of products that are offered increases with the memory length $M$, because it is beneficial to offer more products and avoid repeating the same product within the memory length when satiation effects last for a long period. When the memory length increases, the assortments include more products with lower prices and higher base attraction values, as shown in Figure 2(b) and 2(c).
[…]
Assortment details of four different policies (a)-(d) show the assortments of four polices. The horizontal axes represent products indexed. The vertical axes represent the days of a week. The products offered are colored blue. (e) and (f) show the price and base attraction value of each product.
Figure 5.2.1 presents assortment details for examining how the memory length affects assortment patterns. Figure 5.2.1 (a)-(d) show the assortments obtained from MNL-based models with $M\in\{0,1,2,3\}$. The horizontal axes represent products indexed from the highest to the lowest price. The vertical axes represent the days of a week. The products offered are colored blue. The subfigures show that the interval between two adjacent offerings of the same dish grows when $M$ increases, and optimal assortments represent a cyclic-style structure. The interval for two adjacent offerings of most products is the memory length $M$, which suggests that the cycle length is $M+1$, aligning with the $(M+1)$-cyclic policy in Section 4. Moreover, we observe that products with higher prices or lower base attraction values are more sensitive to the satiation effect and thus will be offered at most once within the span of the memory length. For instance, the high-price products 1 to 5 strictly repeat once in intervals of $M$ days to avoid the satiation effect, possibly because the revenue loss caused by the utility deterioration will be scaled up by the high price. On the other hand, products with relatively high base attraction value may possibly be offered more than once within the memory period, e.g., product 17,18. The price and base attraction value for each product are shown in Figure 5.2.1 (e) and (f) for reference.
@@ 5.2.2 Performance Comparison
In this section, we focus on two metrics to evaluate the assortment performance: variety and revenue, as these are concerns for the corporation cafeteria. We adopt the Herfindahl-Hirschman Index (HHI), a measure of market concentration (Rhoades 1993), to quantify the variety of an assortment planning. The HHI value is $\sum_{i\in[n]}(\frac{h_{i}}{H})^{2}$, where $n$ is the total number of products, $h_{i}$ is the number of times product $i$ is offered, and $H=\sum_{i\in[n]}h_{i}$. A low HHI value reflects a large variety. To compute the out-of-sample revenue, we randomly sample 20 subsets from the second data set. In each subset, we estimate the “ground truth” model with $M=2$ that was selected by cross-validation on the second data set and calculate the revenue based on the “ground truth” demand model. The average revenue across the 20 data sets is used as the out-of-sample revenue to avoid random noise. Besides the assortments discussed in Section 5.2.1, we additionally compute assortments obtained by the $(M+1)$-cyclic policy with non-overlapping constraints that is presented in Section 4.
[…]
Assortment variety and average revenue in the out-of-sample test The horizontal axis represents the Herfindahl-Hirschman Index value of each policy. A small HHI value means a wide variety. The vertical axis shows the average revenue of a policy across 20 test sets. We normalize the revenue of the current practice to one.
Figure 5.2.2 illustrates the relationship between the variety and the revenue of an assortment plan. The horizontal axis represents the HHI value, and the vertical axis shows the average normalized revenue. We normalize the out-of-sample revenue of different assortment plans with the revenue of current practice as a benchmark. Generally, assortments with a low HHI value can help avoid the negative impacts of satiation effects and maintain high product utility. Figure 5.2.2 shows that the assortments generated by our model ($M>0$) have a lower HHI and higher revenue than the assortments obtained with the MNL model. For instance, the average revenue obtained using our model with $M=3$ is approximately $10.4\%$ greater than the revenue under the MNL model. The $(M+1)$-cyclic policy also performs well, with cyclic assortments experiencing only a slight revenue loss compared to the optimal assortments. We also note that the manually decided assortments in practice have low HHI values but carry substantial revenue losses, as in practice, managers may decide the assortments based on many factors that are not observed. These results demonstrate that a wide variety of products with a low frequency can mitigate the satiation effect. Our model achieves this by efficiently balancing products’ utility, revenue, and history-dependent effects.
@@ 6 Numerical Study
This section presents our numerical studies on synthetic data. Section 6.1 illustrates the efficiency of our formulations regarding computation time and optimality gap compared to other formulations. Section 6.2 shows that our formulation generates significantly more revenue than benchmark heuristics. In Section 6.3, we demonstrate that the bound-free model (12) can scale up to solve large-size instances. Our numerical studies were implemented in the Julia (Bezanson et al. 2017), using modeling language JuMP (Lubin et al. 2023), on a Virtual Machine with a 32-core Intel Xeon (Skylake, IBRS) @2.39 GHz processor. All MILP models were solved using Gurobi 11.0.0 (Gurobi Optimization, LLC 2023), and MIECP models were solved using Mosek 10.2 (MOSEK 2024).
@@ 6.1 Formulation Efficiency
We consider four alternate formulations for solving (HAP) without constraints and set the bounds of the no-purchase probability variable as $\rho^{t}_{L}=\frac{1}{({1+\sum_{j\in[N]}\exp(\beta_{j}^{0}+\sum_{m\in I_{j}}\beta_{j}^{m}))}}$ and $\rho^{t}_{U}=1$ for the first three models:
@@
  • $\mathsf{Env}$: This is the mixed-integer linear program obtained by replacing the exponential cone constraint in (Conic) with the convex envelope based constraints (8) and (9), as detailed in Remark 3.3.
  • $\mathsf{ML}$: This is the mixed-integer linear formulation based on a multilinear extension of attraction value functions, and its derivation is detailed in the ecompanion EC.5.1.
  • $\mathsf{SCIP}$: This refers to solving the binary nonlinear model (HAP) by the global optimization solver SCIP.
We consider the satiation effects in this numerical study. For a fixed number of products $N$ and memory length $M$, the base utility of each product $\beta_{i}^{0}$ is drawn from a uniform $[-1,1]$ distribution, and its revenue $r_{i}$ from a uniform $[1,10]$ distribution. The no-purchase utility is 00. For each $i\in[N]$ and $m\in[M]$, we set the history-dependent effect $\beta_{i}^{m}$ from two uniform distributions $U[-1,0]$ and $U[-2,-1]$ to represent a weak and strong satiation effect, denoted as W and S, respectively. We vary $N\in\{30,50,70\}$, $M\in\{1,2\}$, and $T\in\{5,10\}$. For each combination of $N$, $M$, and a satiation level (W or S), we generate five instances. In EC.6, we provide additional computational results on settings where history-dependent effects are mixed, memory lengths are large, and both cross-periods and cross-products constraints are present. To solve our MIECP models with large memory lengths, we propose a cutting-plane algorithm, which is discussed in EC.5.3.
We solve each instance of (HAP) using one of the formulations $\mathsf{Conic}$, $\mathsf{Env}$, $\mathsf{ML}$, and $\mathsf{SCIP}$. We report the computation time and the end gap, which is defined as $G_{\textsf{end}}=100\%\times\frac{R_{\mathsf{U}}-R_{\mathsf{IP}}}{R_{\mathsf{IP}}}$, where $R_{\mathsf{U}}$ and $R_{\mathsf{IP}}$ is the best upper bound and the best integer solution at termination with a time limit of 3600 seconds and an optimality gap tolerance of $0.5\%$, respectively.
Table 2: Computation time and end gap of four formulations under satiation effects
@@
5 1 30 W 3.86 3.22 156.73 - 10 1 30 W 4.38 3.65 179.23 -
5 1 30 S 6.81 5.89 301.08 - 10 1 30 S 7.83 6.56 343.77 -
5 1 50 W 2.89 2.5 295.99 - 10 1 50 W 3.3 2.82 336.43 -
5 1 50 S 5.64 5.04 517.19 - 10 1 50 S 7.11 5.64 588.26 -
5 1 70 W 3.95 3.55 400.53 - 10 1 70 W 5.17 4.77 455.25 -
5 1 70 S 3.84 3.45 713.79 - 10 1 70 S 5.05 4.2 810.67 -
5 2 30 W 4.57 4.57 224.85 - 10 2 30 W 5.64 5.61 275.29 -
5 2 30 S 13.17 13.17 587.57 - 10 2 30 S 15.4 15.32 786.57 -
5 2 50 W 4.77 4.77 430.75 - 10 2 50 W 5.9 5.76 539.85 -
5 2 50 S 10.1 10.06 921.6 - 10 2 50 S 13.34 12.09 1419.53 -
5 2 70 W 5.47 5.3 717.98 - 10 2 70 W 6.94 6.46 940.42 -
5 2 70 S 6.85 6.43 1243.68 - 10 2 70 S 9.02 7.85 1890.09 -
Besides the computation time and end gap, it is also helpful to compare the continuous relaxation of different formulations. We compute the root gap by $100\%\times\frac{R_{\textsf{Rlx}}-R_{\textsf{IP}}}{R_{\textsf{IP}}}$, where $R_{\mathsf{Rlx}}$ is the objective value of the continuous relaxation of a given formulation. A small root gap indicates a tight formulation. Table 3 reports the average root gap of Conic, Env, and ML under different parameter configurations. Overall, the root gap of our formulations Conic and Env is significantly smaller than that of ML. It indicates the tightness of our formulations. Moreover, the root gap of our formulations is stable and does not scale with the problem size. For instance, the root gap of Env is no more than $7\%$ when $M=1$ and no more than $16\%$ when $M=2$.
6.2 Performance of Heuristic Policies
We compare the revenue generated by our formulation against two heuristic policies without constraints.
  • $\mathsf{Sequential-RO}$: The sequential RO-policy shown in Algorithm 1. It can serve as a heuristic policy for settings with satiation effects.
  • $\mathsf{Sequential-LOSPO}$: For each period, this policy drops one satiation product that causes the greatest revenue reduction from the base assortment obtained from the $\mathsf{Sequential-RO}$ policy. We refer to it as sequential leave-one-satiation-product-out (Sequential-LOSPO) policy.
We consider instances with mixed addiction-satiation effects. A product has an addiction (resp. satiation) effect with probability $\theta$ (resp. $1-\theta$), where $\theta\in\{0,0.1,0.2\}$. When $\theta=0$, all products have satiation effects. For a product with addiction (resp. satiation) effect, its $\{\beta_{i}^{m}\}_{m\in[M]}$ are uniformly sampled from $[0,1]$ (resp. $[-2,-1]$). We fix $M=2$ and generate five instances for each combination of $N,T$, and $\theta$.
We compare revenue obtained by the two heuristic policies and our formulation. We define the relative revenue gap as $100\%\times\frac{R_{\mathsf{Env}}-R_{\mathsf{heuristic}}}{R_{\mathsf{Env}}}$, where $R_{\mathsf{heuristic}}$ is revenue generated by one of two heuristic policies. $R_{\mathsf{Env}}$ is computed by $\mathsf{Env}$ within 3600 seconds. Table 4 shows that the revenue gap of both heuristics is not negligible, especially when $\theta$ is small. For instance, the smallest relative gap is at least more than 10% if $\theta=0$, suggesting that our formulation has more advantage in scenarios where the satiation effect is dominant. Although the revenue gap shrinks when $\theta$ increases, our formulation still holds value as it has an optimality guarantee and is easily extended to constrained cases.
[…]
$\mathsf{Sequential-RO}$ $\mathsf{Sequential-LOSPO}$
N T $\theta=0$ $\theta=0.1$ $\theta=0.2$ $\theta=0$ $\theta=0.1$ $\theta=0.2$
30 5 18.60 13.02 2.99 15.09 11.77 2.60
30 10 31.37 19.43 4.18 28.47 18.65 3.81
50 5 14.39 5.18 2.02 12.90 4.63 1.84
50 10 25.72 7.73 2.82 24.20 7.27 2.66
70 5 12.31 4.00 3.59 10.91 3.80 3.40
70 10 21.86 5.74 5.03 20.61 5.59 4.88
6.3 Computing $(M+1)$-Cyclic Policies
@@
In this section, we demonstrate the efficiency and scalability of the bound-free model (12) for computing $(M+1)$-cyclic policies under the non-overlapping constraint. Note that due to the presence of $\boldsymbol{\Gamma}$ variables, the number of variables in (12) is $\mathcal{O}(N^{2})$. To address this issue, we propose a projected-cutting-plane algorithm to obtain a projected approximation of (12). The cutting plane algorithm first sets a base formulation without using $\boldsymbol{\Gamma}$ variables, that is, constraints (12c)-(12e), and solves its continuous relaxation to obtain a continuous solution. Then, the algorithm calls a separation oracle to generate cuts violated by the continuous solution, which are selected from the projection of constraints (12c)-(12e) onto the space of the base formulation, i.e., the space of $(\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{\gamma})$ variables. Adding new cuts to the base model leads to a tighter formulation. Running the separation oracle once is a round of cut generation. We can introduce new cuts over $K$ rounds and refer to the resulting formulation as BF-$K$, where “BF” stands for “bound-free”. We refer to EC.5.2 for a detailed description of the projected cutting-plane algorithm.
Clearly, a larger $K$ leads to a tighter formulation at the cost of adding cuts. To select a proper $K$ that balances the formulation tightness and the formulation size, we compute the optimality gap closed after the $K^{\text{th}}$ round as follows
$\textsf{GClosed}_{K}=100\%\times\frac{\textsf{Opt}_{0}-\textsf{Opt}_{K}}{\textsf{Opt}_{0}-\textsf{Opt}_{\infty}},$
where $\textsf{Opt}_{K}$ is the optimal objective value of the continuous relaxation of BF-$K$, and the index 00 (resp. $\infty$) refers to the base (resp. limiting) model. Table 5(a) reports the gap closed by BF-$K$ for $K\in\{1,2,3\}$ on instances with $M\in\{1,2\}$ and $N\in\{50,100\}$. BF-$1$ closes about $80\%$ of the optimality gap, suggesting setting $K=1$ is good enough.
@@
Table 5: The computational performance of (12) with a projected-cutting-plane implementation
(a) Performance of the cutting-plane algorithm
N M $\textsf{GClosed}_{1}$ $\textsf{GClosed}_{2}$ $\textsf{GClosed}_{3}$
50 1 79.25 91.19 96.21
50 2 82.24 93.14 95.90
100 1 89.82 95.61 97.58
100 2 77.24 90.30 93.48
(b) Performance of bound-free model BF-$1$
N M #$\#\mathsf{sol}$ $T_{\mathsf{opt}\small{(s)}}$ $G_{\mathsf{end}\small{(\%)}}$
50 1 5 0.57 0.30
50 2 5 70.74 0.46
100 1 5 8.19 0.42
100 2 5 378.59 0.49
Table 5(b) reports the performance of BF-$1$. The result demonstrates the scalability and efficiency of the bound-free model compared to (Conic). BF-1 solves all instances to optimality within 400 seconds. In contrast, under the number of products $N=50$ and memory length $M=2$, both Conic and Env fail to solve some of the instances as shown in Table 2.